# Similarity-Distance-Magnitude Activations

### v1

@@ 3.2.1 Controlling Selective Classification Recall
Next, we estimate label-conditional conformal thresholds, $\psi_{1},\ldots\psi_{C}]$, over the output from the $\operatorname{\text{{sdm}}}$ activation among a subset of the distribution determined by progressively larger values of $q^{\prime}$ (among $\left\lfloor{q^{\prime}}\right\rfloor>0$) until all thresholds are at least $\alpha$. By setting our stopping criteria at $\alpha$ rather than $\frac{1}{C}$, we also restrict the region to our empirically-motivated prediction-conditional quantity, $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})_{\hat{y}}$. The procedure appears in Alg. 1. If we find a finite $q^{\prime}_{\rm{min}}$ that obtains such thresholds, we refer to the resulting region as the $\operatorname{\text{{high-reliability}}}$ ($\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$) region, taking membership in this region as our selection criteria:
@@
To calculate this quantity for new, unseen test points ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, we require $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ to calculate $\operatorname{\text{$q$}}$ and $\operatorname{\text{$d_{\rm{nearest}}$}}$; the cached class-wise empirical CDFs of the distances over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ of Eq. 5; and $q^{\prime}_{\rm{min}}$ and the thresholds ($\psi_{1},\ldots\psi_{C}]$). Evaluation of the $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ selection criteria is straightforward: We simply assess the conditional accuracies for the admitted points after stratifying by the predictions and the true labels, each in turn.
@@
Algorithm 1 Search Algorithm to Find $q^{\prime}_{\rm{min}}$ to Determine the $\operatorname{\text{{high-reliability}}}$ Region
1:cached $(q^{\prime},\operatorname{\text{{sdm}}}({\bm{z}}^{\prime}))$ for $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, $\alpha\in(\frac{1}{C},1]$
2:procedure find-min-rescaled-similarity(cached $(q^{\prime},\operatorname{\text{{sdm}}}({\bm{z}}^{\prime}))$ for $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, $\alpha\in(\frac{1}{C},1]$)
@@
4:  $\psi_{1},\ldots\psi_{C}]\leftarrow[\infty,\ldots,\infty]$
@@
8:    Calculate $\psi_{c}={\rm{inverseCDF}}^{y_{c}}_{\rm{ca}}(1-\alpha^{\prime})~\forall~c~\in\{1,\ldots,C\}$ $\triangleright$ Quantile functions are inverses of L. 7
@@ 4 Experiments
We provide controlled comparisons of our proposed methods over representative LLMs and tasks, systematically ablating relevant components, holding the data and underlying LLM constant, ceteris paribus. We consider in-distribution, co-variate shifted, and out-of-distribution test sets. We consider representative estimators over the existing LLM architecture (i.e., without additional parameters); with CNN adaptors; and with the $\operatorname{\text{{sdm}}}$ activation layer.
@@ 4.1 Task: Binary Sentiment Classification
Our first task is predicting the sentiment of movie reviews. We use the commonly used benchmark data of Maas et al. (2011). This is a binary classification task with $y\in\{0={\rm{negative}},1={\rm{positive}}\}$. $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ are constructed from a total of 18k instances. The held-out set for evaluation, $|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|=1583$, is from the same distribution as $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$. This is a well-studied task for which the surface-level signals correlated with the target labels are expected to be effectively modeled by large parameter LLMs; as such, relatively high task accuracies are expected. We use the label $\operatorname{\text{{Sentiment}}}$ for the in-distribution test set.
@@
In the Appendix, we consider two additional out-of-distribution challenge test sets, $\operatorname{\text{{SentimentShuffled}}}$ and $\operatorname{\text{{SentimentOODShuffled}}}$, constructed by randomly shuffling the input documents for each of $\operatorname{\text{{Sentiment}}}$ and $\operatorname{\text{{SentimentOOD}}}$, respectively. The semantics of the original labels are maintained. This requires the models and estimators to attempt a sentiment classification over the bag-of-words input, or reject the classification. This represents the setting where an LLM is given far out-of-distribution input, and additionally provides a control on test-set contamination of the underlying LLMs, which due to the shuffling, are relatively unlikely to have seen all of the long, contiguous n-gram sequences from these documents in training or fine-tuning.
@@
For this task, we prompt the LLMs for a binary classification (Yes or No) as follows: Here is a movie review. <review> DOCUMENT </review> Is the sentiment of the movie review positive? Answer Yes if the sentiment is positive. Answer No if the sentiment is negative. Start your response with Yes or No.
@@ 4.2 Task: $\operatorname{\text{{Factcheck}}}$
As a more challenging binary classification task for LLMs, we consider the fact check data of Azaria & Mitchell (2023). The training and calibration sets, a combined total of 6k instances, consist of single sentence statements that have been semi-automatically generated via templates and a knowledge base. The task is to determine whether the statement is true or false, $y\in\{0={\rm{false}},1={\rm{true}}\}$. The held-out eval set ($\operatorname{\text{{Factcheck}}}$), $|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|=245$, the focus of our analysis, has been constructed by having an LLM generate a statement continued from a true statement not otherwise in the dataset. These evaluation statements are checked manually and assigned labels by human annotators. In addition to being a relatively challenging task that evaluates—at least in principle—the latent knowledge stored within an LLM’s parameters, the test set is representative of the types of co-variate shifts over high-dimensional inputs that can be problematic for real applications, and challenging to characterize without model assistance and ground-truth labels. It was observed in Azaria & Mitchell (2023) that the accuracy of existing LLM classifiers is lower on this generated, held-out test set compared to the calibration set. However, these test sentences would seem to also be simple true-false statements, reflecting that it is not always immediately obvious for a human user to detect distribution shifts over high-dimensional inputs. As such, we seek for our models and estimators to reflect such shifts via the predictive uncertainty.
@@
Similar to the sentiment task, we prompt the LLMs for a binary classification (Yes or No) as follows: Here is a statement that may contain errors. <statement> DOCUMENT </statement> Is the statement true? Answer Yes if the statement is true. Answer No if the statement is false. Start your response with Yes or No.
@@ 4.3 Models
Hidden-states.
For both models, we take as $\bm{h}$} the concatenation of the final-layer hidden state of the final sequence position (i.e., the hidden state that is the input to the linear-layer over the output vocabulary for the Yes or No generation) with the sequence-wise mean over all final-layer hidden states. For $\operatorname{\text{{phi3.5}}}$, this results in ${\bm{h}}\in\mathbb{R}^{6144}$, and for $\operatorname{\text{{Mixtral8x7B}}}$, ${\bm{h}}\in\mathbb{R}^{8192}$.
@@ 4.4 Estimators
Holding each underlying LLM constant, we examine representative classes of estimators used with neural networks, setting $\alpha=0.95$ for all experiments. At the most basic, but also, perhaps the most commonly used in practice, representing the absence of a post-hoc calibration method, we simply threshold the output, $\operatorname{softmax}({\bm{z}})\geq\alpha$, where the temperature $\tau=1$. As an established empirical approach for calibrating neural networks, we provide a comparison to temperature scaling (Guo et al., 2017), a single parameter version of post-hoc Platt-scaling (Platt, 1999), with the label $\operatorname{\text{{tempScaling}}}$. In this case, the estimator is the thresholding of the output $\operatorname{softmax}({\bm{z}};\tau)\geq\alpha$ after learning a value for $\tau$ over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$. We also provide a comparison to two representative conformal predictors, the $\operatorname{\textsc{APS}}$ method of Romano et al. (2020) and the adaptiveness-optimized $\operatorname{\textsc{RAPS}}$ algorithm of Angelopoulos et al. (2021). The admission criteria for the $\operatorname{\textsc{APS}}$ and $\operatorname{\textsc{RAPS}}$ estimators is prediction sets of size 1, at the $0.05$ level (i.e., $1-\alpha$, as defined here). We consider these estimators over the logits corresponding to the Yes and No indexes of the output linear-layer of the underlying LLM ($\operatorname{\text{{phi3.5}}}$ and $\operatorname{\text{{Mixtral8x7B}}}$), which provides a reference point without introducing additional adaptor layers. We also consider these baselines over 1-D CNN adaptors over $\bm{h}$} of each LLM ($\operatorname{\text{{phi3.5+adaptor}}}$ and $\operatorname{\text{{Mixtral8x7B+adaptor}}}$). These baseline adaptors are identical to those used in the corresponding $\operatorname{\text{{sdm}}}$ activation layers, with $M=1000$, and are similarly trained for $J=10$ iterations of 200 epochs, but unlike the $\operatorname{\text{{sdm}}}$ activation layers, the stopping criteria is the maximum balanced (across classes) average cross-entropy loss.
@@
We then compare to the final-layer $\operatorname{\text{{sdm}}}$ activations over $\bm{h}$} of each LLM ($\operatorname{\text{{phi3.5+sdm}}}$ and $\operatorname{\text{{Mixtral8x7B+sdm}}}$). For reference, we provide the result of thresholding a $\operatorname{softmax}$ over these adaptors at $\alpha$, as above, as well as a thresholded $\operatorname{softmax}$ that simply treats $\operatorname{\text{$d$}}$ as the inverse-temperature, $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$, which is equivalent to setting q=e2q=e-2 in the $\operatorname{\text{{sdm}}}$ activation. We also consider an analogous threshold over the activation output, $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})\geq\alpha$, for which we use $\operatorname{\text{{sdm}}}$ as the estimator label. Finally, we use the label $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ for the selection criteria proposed in Eq. 9.
@@ 5 Results
The results for the in-distribution $\operatorname{\text{{Sentiment}}}$ dataset appear in Table 1. Even on this in-distribution dataset, the estimators over the underlying LLMs without adaptor layers exhibit modest over-confidence, which is reflected in conditional accuracies that fall below the expected $\alpha$. The estimators over the adaptor layers and the $\operatorname{\text{{sdm}}}$ activation all obtain the desired conditional accuracies, with the class-wise accuracies of the models themselves $\geq\alpha$, with differences arising in the proportion of admitted points. Here and elsewhere, $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ tends to be overly conservative in rejecting points. To be expected, the $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ estimator tends to be more conservative than simply thresholding the $\operatorname{\text{{sdm}}}$ activation on the in-distribution data, but the latter lacks the assurances on the recall obtained by the constraints on the $\operatorname{\text{{high-reliability}}}$ region. In practice, this behavior can be used as a basis to triage selective classifications: For example, documents in the $\operatorname{\text{{high-reliability}}}$ region might be treated as automated, or semi-automated, predictions in the decision pipeline, whereas other documents might be triaged by $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})_{\hat{y}}$ for calling more resource-intensive LLM tools, or human adjudication. Importantly, as we discuss below with the co-variate-shifted and out-of-distribution datasets, just using the $\operatorname{softmax}$ or the other estimators does not provide a reliable substrate for basing such conditional branching decisions.
@@
Co-variate-shifted and Out-of-distribution data.
Understanding $q^{\prime}_{\rm{min}}$.
For reference, Table 3 provides the results over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ for the $\operatorname{\text{{sdm}}}$-based estimates. The value of $q^{\prime}_{\rm{min}}$ tends to increase as the accuracy over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ decreases, reflecting a more conservative $\operatorname{\text{{high-reliability}}}$ region that admits fewer points. Alg. 1 failed to find a finite $q^{\prime}_{\rm{min}}$ for $\operatorname{\text{{Mixtral8x7B+sdm}}}$ over the $\operatorname{\text{{Factcheck}}}$ calibration set at $\alpha=0.95$, so for reference, we also show the $\operatorname{\text{{high-reliability}}}$ region at $\alpha=0.94$, as well as in Table 2. In this way, $q^{\prime}_{\rm{min}}$ provides a principled, data-driven indicator of the reliability of the estimates, which is interpretable as a simple indicator as to whether the conditional accuracies are, or are not, obtainable over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ at the specified $\alpha$.
Far out-of-distribution data.

### v2

@@ 3.2.1 Controlling the Class-conditional Accuracy among Selective Classifications with SDM Estimators
Next, we estimate label-conditional conformal thresholds, $\psi_{1},\ldots\psi_{C}]$, over the output from the $\operatorname{\text{{sdm}}}$ activation among a subset of the distribution constrained by progressively larger values of $q^{\prime}$ (among $\left\lfloor{q^{\prime}}\right\rfloor>0$) until all thresholds are at least $\alpha$. By setting our stopping criteria at $\alpha$ rather than $\frac{1}{C}$, we also restrict the region to our empirically-motivated prediction-conditional quantity, $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})_{\hat{y}}$. The procedure appears in Alg. 1. If we find a finite $q^{\prime}_{\rm{min}}$ that obtains such thresholds, we refer to the resulting region as the $\operatorname{\text{{high-reliability}}}$ ($\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$) region, taking membership in this region as our selection criteria:
@@
To calculate this quantity for new, unseen test points ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, we require $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ to calculate $\operatorname{\text{$q$}}$ and $\operatorname{\text{$d_{\rm{nearest}}$}}$; the cached class-wise empirical CDFs of the distances over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ of Eq. 5; and $q^{\prime}_{\rm{min}}$ and the thresholds, $\psi_{1},\ldots\psi_{C}]$. Evaluation of the $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ selection criteria is straightforward: We simply assess the conditional accuracies for the admitted points after stratifying by the predictions and the true labels, each in turn.
@@
Algorithm 1 Search Algorithm to Find $q^{\prime}_{\rm{min}}$ and $\psi_{1},\ldots\psi_{C}]$ to Estimate the $\operatorname{\text{{high-reliability}}}$ Region
1:cached $(q^{\prime},\operatorname{\text{{sdm}}}({\bm{z}}^{\prime}))$ $\forall~{\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, $\alpha\in(\frac{1}{C},1]$
2:procedure estimate-high-reliability-region(cached $(q^{\prime},\operatorname{\text{{sdm}}}({\bm{z}}^{\prime}))$ $\forall~{\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, $\alpha\in(\frac{1}{C},1]$)
@@
4:  $\psi_{1},\ldots\psi_{C}]\leftarrow[\infty,\ldots,\infty] \triangleright$ Class-wise output thresholds
@@
8:    Calculate $\psi_{c}={\rm{inverseCDF}}^{y_{c}}_{\rm{ca}}(1-\alpha)~\forall~c~\in\{1,\ldots,C\}$ $\triangleright$ Quantile functions are inverses of L. 7
@@
The convention is to refer to the basic architecture of Eq. 6 as the $\operatorname{\text{{sdm}}}$ activation function, and using the activation with the selection criteria of Eq. 9 as the $\operatorname{\text{{sdm}}}$ estimator.
@@ 4 Experiments
We provide controlled comparisons of our proposed methods over representative LMs and tasks, systematically ablating relevant components, holding the data and underlying LM constant, ceteris paribus. We consider in-distribution, co-variate shifted, and out-of-distribution test sets. We consider representative estimators over the existing LM architecture (i.e., without additional parameters); with CNN adaptors; and with the $\operatorname{\text{{sdm}}}$ activation layer.
@@ 4.1 Task: Binary Sentiment Classification
Our first task is predicting the sentiment of movie reviews. We use the commonly used benchmark data of Maas et al. (2011). This is a binary classification task with $y\in\{0={\rm{negative}},1={\rm{positive}}\}$. $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ are constructed from a total of 18k instances. The held-out set for evaluation, $|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|=1583$, is from the same distribution as $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$. This is a well-studied task for which the surface-level signals correlated with the target labels are expected to be effectively modeled by large parameter LMs; as such, relatively high task accuracies are expected. We use the label $\operatorname{\text{{Sentiment}}}$ for the in-distribution test set.
@@
In the Appendix, we consider two additional out-of-distribution challenge test sets, $\operatorname{\text{{SentimentShuffled}}}$ and $\operatorname{\text{{SentimentOODShuffled}}}$, constructed by randomly shuffling the input documents for each of $\operatorname{\text{{Sentiment}}}$ and $\operatorname{\text{{SentimentOOD}}}$, respectively. The semantics of the original labels are maintained. This requires the models and estimators to attempt a sentiment classification over the bag-of-words input, or reject the classification. This represents the setting where an LM is given far out-of-distribution input, and additionally provides a control on test-set contamination of the underlying LMs, which due to the shuffling, are relatively unlikely to have seen all of the long, contiguous n-gram sequences from these documents in training or fine-tuning.
@@
For this task, we prompt the LMs for a binary classification (Yes or No) as follows: Here is a movie review. <review> DOCUMENT </review> Is the sentiment of the movie review positive? Answer Yes if the sentiment is positive. Answer No if the sentiment is negative. Start your response with Yes or No.
@@ 4.2 Task: $\operatorname{\text{{Factcheck}}}$
As a more challenging binary classification task for LMs, we consider the fact check data of Azaria & Mitchell (2023). The training and calibration sets, a combined total of 6k instances, consist of single sentence statements that have been semi-automatically generated via templates and a knowledge base. The task is to determine whether the statement is true or false, $y\in\{0={\rm{false}},1={\rm{true}}\}$. The held-out eval set ($\operatorname{\text{{Factcheck}}}$), $|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|=245$, the focus of our analysis, has been constructed by having an LM generate a statement continued from a true statement not otherwise in the dataset. These evaluation statements are checked manually and assigned labels by human annotators. In addition to being a relatively challenging task that evaluates—at least in principle—the latent knowledge stored within an LM’s parameters, the test set is representative of the types of co-variate shifts over high-dimensional inputs that can be problematic for real applications, and challenging to characterize without model assistance and ground-truth labels. It was observed in Azaria & Mitchell (2023) that the accuracy of existing LM classifiers is lower on this generated, held-out test set compared to the calibration set. However, these test sentences would seem to also be simple true-false statements, reflecting that it is not necessarily straightforward for a human user to detect distribution shifts over high-dimensional inputs. As such, we seek for our models and estimators to reflect such shifts via the predictive uncertainty.
@@
Similar to the sentiment task, we prompt the LMs for a binary classification (Yes or No) as follows: Here is a statement that may contain errors. <statement> DOCUMENT </statement> Is the statement true? Answer Yes if the statement is true. Answer No if the statement is false. Start your response with Yes or No.
@@ 4.3 Models
Hidden states.
For both models, we take as $\bm{h}$} the concatenation of the final-layer hidden state of the final sequence position (i.e., the hidden state that is the input to the linear-layer over the output vocabulary for the Yes or No generation) with the mean over all final-layer hidden states. For $\operatorname{\text{{phi3.5}}}$, this results in ${\bm{h}}\in\mathbb{R}^{6144}$, and for $\operatorname{\text{{Mixtral8x7B}}}$, ${\bm{h}}\in\mathbb{R}^{8192}$.
@@ 4.4 Estimators
Holding each underlying LM constant, we examine representative classes of estimators used with neural networks, setting $\alpha=0.95$ for all experiments. At the most basic, but also perhaps the most commonly used in practice, representing the absence of a post-hoc calibration method, we simply threshold the output, $\operatorname{softmax}({\bm{z}})\geq\alpha$, where the temperature $\tau=1$. As an established empirical approach for calibrating neural networks, we provide a comparison to temperature scaling (Guo et al., 2017), a single parameter version of post-hoc Platt-scaling (Platt, 1999), with the label $\operatorname{\text{{tempScaling}}}$. In this case, the estimator is the thresholding of the output $\operatorname{softmax}({\bm{z}};\tau)\geq\alpha$ after learning a value for $\tau$ over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$. We also provide a comparison to two representative conformal predictors, the $\operatorname{\textsc{APS}}$ method of Romano et al. (2020) and the adaptiveness-optimized $\operatorname{\textsc{RAPS}}$ algorithm of Angelopoulos et al. (2021). The admission criteria for the $\operatorname{\textsc{APS}}$ and $\operatorname{\textsc{RAPS}}$ estimators is prediction sets of size 1, at the $0.05$ level (i.e., $1-\alpha$, as defined here). We consider these estimators over the logits corresponding to the Yes and No indexes of the output linear-layer of the underlying LM ($\operatorname{\text{{phi3.5}}}$ and $\operatorname{\text{{Mixtral8x7B}}}$), which provides a reference point without introducing additional adaptor layers. We also consider these baselines over 1-D CNN adaptors over $\bm{h}$} of each LM ($\operatorname{\text{{phi3.5+adaptor}}}$ and $\operatorname{\text{{Mixtral8x7B+adaptor}}}$). These baseline adaptors are identical to those used in the corresponding $\operatorname{\text{{sdm}}}$ activation layers, with $M=1000$, and are similarly trained for $J=10$ iterations of 200 epochs, but unlike the $\operatorname{\text{{sdm}}}$ activation layers, the stopping criteria is the minimum balanced (across classes) average cross-entropy loss.
@@
We then compare to the final-layer $\operatorname{\text{{sdm}}}$ activations over $\bm{h}$} of each LM ($\operatorname{\text{{phi3.5+sdm}}}$ and $\operatorname{\text{{Mixtral8x7B+sdm}}}$). For reference, we provide the result of thresholding a $\operatorname{softmax}$ over these adaptors at $\alpha$, as above, as well as a thresholded $\operatorname{softmax}$ that simply treats $\operatorname{\text{$d$}}$ as the inverse-temperature, $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$, which is equivalent to setting q=e2q=e-2 in the $\operatorname{\text{{sdm}}}$ activation. We also consider an analogous threshold over the activation output, $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})\geq\alpha$, for which we use $\operatorname{\text{{sdm}}}$ as the estimator label. Finally, we use the label $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ for the selection criteria proposed in Eq. 9.
@@ 5 Results
The results for the in-distribution $\operatorname{\text{{Sentiment}}}$ dataset appear in Table 1. Even on this in-distribution dataset, the estimators over the underlying LMs without adaptor layers exhibit modest over-confidence, which is reflected in conditional accuracies that fall below the expected $\alpha$. The estimators over the adaptor layers and the $\operatorname{\text{{sdm}}}$ activation all obtain the desired conditional accuracies, with the class-wise accuracies of the models themselves $\geq\alpha$, with differences arising in the proportion of admitted points. Here and elsewhere, $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ tends to be overly conservative in rejecting points. To be expected, the $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ estimator tends to be more conservative than simply thresholding the $\operatorname{\text{{sdm}}}$ activation on the in-distribution data, but the latter lacks the assurances on the class-conditional accuracy obtained by the constraints on the $\operatorname{\text{{high-reliability}}}$ region. In practice, this behavior can be used as a basis to triage selective classifications: For example, documents in the $\operatorname{\text{{high-reliability}}}$ region might be treated as automated, or semi-automated, predictions in the decision pipeline, whereas other documents might be triaged by $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})_{\hat{y}}$ for calling more resource-intensive LM tools, or human adjudication. Importantly, as we discuss below with the co-variate-shifted and out-of-distribution datasets, just using the $\operatorname{softmax}$ or the other estimators does not provide a reliable substrate for basing such conditional branching decisions.
@@
Co-variate-shifted and Out-of-distribution data.
Understanding $q^{\prime}_{\rm{min}}$.
For reference, Table 3 provides the results over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ for the $\operatorname{\text{{sdm}}}$-based estimators. The value of $q^{\prime}_{\rm{min}}$ tends to increase as the accuracy over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ decreases, reflecting a more conservative $\operatorname{\text{{high-reliability}}}$ region that admits fewer points. Alg. 1 failed to find a finite $q^{\prime}_{\rm{min}}$ for $\operatorname{\text{{Mixtral8x7B+sdm}}}$ over the $\operatorname{\text{{Factcheck}}}$ calibration set at $\alpha=0.95$, so for reference, we also show the $\operatorname{\text{{high-reliability}}}$ region at $\alpha=0.94$, as well as in Table 2. In this way, $q^{\prime}_{\rm{min}}$ provides a principled, data-driven indicator of the reliability of the estimates, which is interpretable as a simple indicator as to whether the conditional accuracies are, or are not, obtainable over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ at the specified $\alpha$.
Far out-of-distribution data.